Determine the empirical and molecular formula given the percent compositions or masses.


20.2% Sodium, 37.6%Sulfur, 42.2% Oxygen, Molar mass: 682.8g/mol

1
Expert's answer
2021-11-18T12:28:02-0500

20.2/23=0.88

37.6/32=1.175

42.2/16=2.6375


0.88/1.175/2.6375=3/4/9

M(Na3S4O9)=23*3+32*4+16*9=341g/mol

2*341=682

Na6S8O18